I have tried a lot of times, but still cannot figure out how to solve these problems. Please help. And please submit the calcultions too if necessary, because I think I've tried right methods and can't spot where I have mistaken.


1. Solve the equations uusing Cramer's rule.


 1/x + 1/y + 1/z = 1, 2/x + 5/y + 3/z = 0, 1/x + 2/y + 4/z = 3  


2. The vertex of a parabola is (-2,3) and its directrix is x+7=0; find the quation of the parabola.

 

Grade:12

1 Answers

Vikas TU
10 years ago

1) 

1/x + 1/y + 1/z = 1, 2/x + 5/y + 3/z = 0, 1/x + 2/y + 4/z = 3 

Thus, it is a non homogenous eqn.

1. first step calculate |D| i.e determinant.

 

D = 17/120 which is not equals to 0 thus it will have unique solution.
and the solutions will be
x = Dx/D

y = Dy/D

z = Dz/D

Calculate thus and get the ansr.

 

PLZ APPROVE!